# Elliptic butterflies

### v1

@@ 10. Reduction
Under the hypotheses at the beginning of Section 5 and assuming that the base field $\mathbf{K}$} has odd characteristic, we let $b$ be a point on $E$ defined over a separable closure ${\mathbf{K}}_{s}$} of $\mathbf{K}$}. We assume that $db\not=0$. We call $B$ the coset of $b$ under the action of $\langle t\rangle$}. So
@@
We assume that $B$ is left invariant by the absolute Galois group of $\mathbf{K}$}. For every $l$ in ${\mathbf{Z}}/d{\mathbf{Z}}$ we denote by $x_{l}$ the function defined by
@@
There exists a constant $\mathcal{Q}$} such that the following is true. Let $\mathbf{K}$} be a field with odd characteristic. Let $E$ be an elliptic curve over $\mathbf{K}$}. Let $\delta\geqslant 1$ be an integer. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. Let $b$ be a point on $E$ defined over a separable extension of $\mathbf{K}$}. Assume $db\not=0$. Let
@@
This results from the discussion above. To make the procedure clearer, we present the reduction routine as brief pseudo-code in Figure 4.
$\square$
@@
Parameters The constants $\vec{{\mathfrak{a}}}\,^{\prime}$ is defined by Equation (8) and $\vec{{\mathfrak{a}}}\,^{1}$ denotes $\vec{{\mathfrak{a}}}\,_{[\,d^{\prime},\ d-1\,]}-\vec{{\mathfrak{a}}}\,_{[\,0,\ d^{\prime}-1\,]}$ . The constants $\vec{{\mathfrak{e}}}$, $\vec{{\mathfrak{f}}}$ and $\vec{{\mathfrak{i}}}$ are defined by Equation (33). The constant $\vec{{\mathfrak{h}}}$ is defined by Equation (29). The constant $\vec{{\mathfrak{l}}}$ is defined by Equation (32), restricted to $l\in[0,\,d^{\prime}-1]$ and $\mathfrak{l}$$_{\ast}$ denotes the constant ${\mathfrak{l}}_{d^{\prime}}$. The constant $\vec{{\mathfrak{v}}}\,^{\prime}$ denotes $(v^{\prime}_{l}(U^{\prime}))_{l\in{\mathbf{Z}}/d^{\prime}{\mathbf{Z}}}$ . The constant $\vec{{\mathfrak{p}}}$ denotes the vector $\vec{{\mathfrak{e}}}\,\diamond\,\vec{{\mathfrak{v}}}\,^{\prime}+\vec{{\mathfrak{f}}}\,\diamond\,(x^{\prime}(U^{\prime}+l\,t^{\prime}))_{l\in{\mathbf{Z}}/d^{\prime}{\mathbf{Z}}}$ , and $\mathfrak{p}$$_{\ast}$ denotes the constant ${\mathfrak{p}}_{0}-{\mathfrak{f}}_{0}\,x^{\prime}(U^{\prime})-y^{\prime}(U^{\prime})$ .
@@ 11. Multiplication in the residue ring of a fiber
Under the hypotheses and notation of Section 10 we explain how to multiply in the residue ring $\mathbf{L}$} at $B$. We assume that there exists a $\mathbf{K}$}-rational point $R$ on $E$ such that $dR\not=0$. The residue ring $\mathbf{L}$} at $B$ is an algebra of dimension $d$ over $\mathbf{K}$}. We let $(u_{l})_{l\in{\mathbf{Z}}/d{\mathbf{Z}}}$ be the functions on $E$ defined in Section 5. For every $l$ in ${\mathbf{Z}}/d{\mathbf{Z}}$ we let
@@
in $\mathbf{L}$}, given by their coordinates in the basis $\Theta$. We follow [CL09, Section 4.3.4.]. We first define two functions
@@
This finishes the proof of Proposition 6 below. One can summarize this proof by saying that the equivariant version of Chudnovsky’s multiplication algorithm summarized in [CL09, Lemma 6] involves convolutions by constant vectors corresponding to evaluation, interpolation and reduction. And the main consequence of the recursive approach presented in Sections 7, 8, and 10 is that these convolution products are achieved in time ${\mathcal{O}}d\log(d)$ when $d$ is a power of two.
@@
Description function NormalBasis_Multiply($\vec{f},\vec{g}$)   $\vec{\alpha}\leftarrow\textsc{Butterfly\_Evaluate}\,(\vec{f})$   $\vec{\beta}\leftarrow\textsc{Butterfly\_Evaluate}\,(\vec{g})$   $\vec{H}\leftarrow(\vec{f}-\sigma(\vec{f}))\,\diamond\,(\vec{g}-\sigma(\vec{g}))$   $\vec{h}\leftarrow\textsc{Butterfly\_Reduce}_{R}\,(\vec{H})$ $\triangleright$ Reduction modulo $R+{\langle t\rangle}$   $\vec{\gamma}\leftarrow\textsc{Butterfly\_Evaluate}\,(\vec{h})$   $\vec{\delta}\leftarrow\vec{\alpha}\ \diamond\ \vec{\beta}-\vec{\gamma}$   $\vec{k}\leftarrow\textsc{Butterfly\_Interpolate}\,(\vec{\delta})$   $\vec{h}\leftarrow\textsc{Butterfly\_Reduce}_{B}\,(\vec{H})$ $\triangleright$ Reduction modulo $b+{\langle t\rangle}$   return $\vec{h}+\vec{k}$ end function
Parameters The function $\textsc{Butterfly\_Reduce}_{R}$ (resp. $\textsc{Butterfly\_Reduce}_{B}$) denotes the reduction routine defined in Figure 4, where the constants $\mathfrak{h}$} and $\mathfrak{l}$} depend on the point $R\in E({\mathbf{K}})$ (resp. $b\in E({{\mathbf{K}}_{s}})$) as defined in Section 13.
@@
There exists a constant $\mathcal{Q}$} such that the following is true. Let $\mathbf{K}$} be a field with odd characteristic. Let $E$ be an elliptic curve over $\mathbf{K}$}. Let $\delta\geqslant 1$ be an integer. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. We assume that there exists a point $R$ in $E({\mathbf{K}})$ such that $dR\not=0$. Let $b$ be a point on $E$ defined over a separable extension of $\mathbf{K}$}. Assume $db\not=0$. Let
@@ 12. Implementation and experimental results
A public implementation of the algorithms presented in this work, developed in the computational algebra system Magma [BCP97], is available online [CL25]. In particular, it includes the classical Fast Fourier Transform (FFT) with the Cooley–Tukey algorithm, as well as our routines, both requiring only $O(d\log(d))$ operations in $\mathbb{F}_{q}$ for evaluations or interpolations at $d$ points.
We present in Table 1 the timings measured for our Magma routines on a standard laptop, working modulo a 64-bit prime. These results are provided for comparison purposes only, given the likely overhead of the Magma interpreter in these highly recursive routines. The relative differences in timings, namely, a constant factor of about $5$ between Cooley-Tukey and the elliptic butterflies, are nonetheless significant in practice.
$\log_{2}(d)$ Cooley-Tukey [Wik] Elliptic Butterflies
Evaluate Interpolate Evaluate (Fig. 2) Interpolate (Fig. 3)
8 0.00 s 0.01 s 0.00 s 0.01 s
9 0.01 s 0.01 s 0.02 s 0.01 s
10 0.02 s 0.03 s 0.03 s 0.03 s
11 0.04 s 0.03 s 0.06 s 0.07 s
12 0.03 s 0.03 s 0.13 s 0.15 s
13 0.05 s 0.06 s 0.28 s 0.31 s
14 0.12 s 0.12 s 0.56 s 0.66 s
15 0.25 s 0.25 s 1.17 s 1.38 s
16 0.54 s 0.55 s 2.45 s 2.72 s
@@ 13. Elliptic bases for finite field extensions
We have shown in [CL09, Theorem 2] how evaluation and interpolation of elliptic functions provide normal bases for finite field extensions having quasi-linear time multiplication. We consider in this section extensions of degree $d=2^{\delta}$, a power of two. We construct normal bases allowing multiplication in time ${\mathcal{Q}}d\log(d)$ for these extensions. We prove the theorem below.
@@
We now consider the case $\nu\leqslant\delta$ and use the constructions in [CL09]. The integer $d_{q}$ as defined in [CL09, Definition 1] is either $d^{2}$ or $2d^{2}$. Since $4d^{4}\leqslant q$ we have $d_{q}\leqslant\sqrt{q}$ and we can apply [CL09, Lemma 9]. There exists an elliptic curve $E$ over $\mathbf{K}$}, a point $t$ of order $d$ in $E({\mathbf{K}})$, a point $R$ in $E({\mathbf{K}})$ such that $dR\not=0$, and a point $b$ in $E({{\mathbf{K}}_{s}})$ such that $db\not=0$ and the conjugate of $b$ by the Frobenius of $E/{\mathbf{K}}$ is $b+t$. We set $B=b+{\langle t\rangle}$. This is a degree $d$ divisor on $E$. It is irreducible over $\mathbf{K}$}. The residue field at $B$ is a degree $d$ extension field of $\mathbf{K}$} which we denote by $\mathbf{L}$}. The $\theta_{l}$ introduced in Equation 34 form a normal basis $\Theta$ of $\mathbf{L}$}. According to Proposition 6 there exists a straight-line program that multiplies to elements in $\mathbf{L}$} given by their coordinates in the basis $\Theta$ at the expense of ${\mathcal{Q}}d\log(d)$ operations in $\mathbf{K}$}.
$\square$
@@ 14. Elliptic Goppa codes
The length of the Hasse interval is $4\sqrt{q}$. So there are two consecutive multiples of $d^{2}$ in it. At least one of them is not congruent to $1$ modulo $p$. We deduce that there exists an elliptic curve $E$ over $\mathbf{K}$} such that $E({\mathbf{K}})$ contains a point $t$ of order $d$. We set
@@
If the check is failed we deduce that there are errors. We then remind that $C$ is a genus 0 code under its clothes of genus $1$ code. In this context, decoding up to half the minimum distance is achieved thanks to an half-gcd computation as explained in [Sug+75]. The time complexity is $\log(d)$ times the complexity of multiplying two polynomials of degree $d$ and coefficients in $\mathbf{K}$}. See [GG13, Chapter 11]. The following theorem summarizes the content of this section.
@@ 15. Elliptic LWE cryptography
Our third application concerns the construction of secure cryptographic schemes within quantum computational models, and in particular in the presence of quantum algorithms such as those of Shor [Sho97] and Grover [Gro96]. Among the cryptosystems most extensively studied in this context are those based on computational problems over Euclidean lattices.
@@
Definition 10 (Short-LWE assumptions [Reg05], [Reg09, Lemma 4.4]).
Let $d,q\geq 2$ be integers, let $\chi$ be an error distribution over $\mathbf{Z}$} (typically a discrete Gaussian), and let $\vec{s}\leftarrow\chi^{m}$ be a secret vector. Given pairs $(\mathsf{A},\mathsf{A}\vec{s}+\vec{e})$, where $\mathsf{A}\leftarrow({\mathbf{Z}}/q{\mathbf{Z}})^{m\times d}$ is uniformly random and $\vec{e}\leftarrow\chi^{m}$,
@@
Let $E$ be an elliptic curve defined modulo a prime $q$, with a point $t\in E(\mathbb{Z}/q\mathbb{Z})$ of order $d=2^{\delta}$, and let $b$ be another point in $E(\mathbb{Z}/q\mathbb{Z})$ such that $db\neq 0$. Let $\otimes$ denote the multiplication law modulo $b+\langle t\rangle$ defined in Proposition 6. Finally, let $\chi$ be an error distribution over $\mathbb{Z}$ (typically a discrete Gaussian), and let $\vec{s}\leftarrow\chi^{d}$ be a secret vector.
@@
Transposing the Regev construction in this setting yields a CPA-secure encryption scheme.
@@
The correctness follows from the equation $p=\langle\vec{r}\,,\vec{e}\rangle-\langle\vec{e}_{1}\,,\vec{s}\rangle+e_{2}+\mu\,\lfloor q/2\rceil\,.$ Since the error term remains small for appropriate parameter choices, decryption succeeds with high probability. Security in the chosen-plaintext attack (CPA) model follows directly from the decisional LWE assumption. The vector $\vec{w}$ is indistinguishable from a random vector, which makes $\vec{c}_{1}$ and $\langle\vec{r},\,\vec{w}\rangle+e_{2}$ two LWE instances, each indistinguishable from uniform, thereby perfectly hiding the message.
In terms of complexity, applying $\phi_{\vec{a}}$ requires $\mathcal{Q}\,d\log(d)$ operations in $\mathbb{Z}/q\mathbb{Z}$, using the algorithms described in Section 10. By Tellegen’s principle, applying the transpose map $^{\mathrm{t}}\phi_{\vec{a}}$ has the same asymptotic complexity [KKB88, BLS03]. More generally, all steps can be carried out within this time bound. The sizes of the keys and ciphertexts are $O(d)$ bits.
Parameters Let $E$ be an elliptic curve defined modulo a prime $q$ with a point $t\in E(\mathbb{Z}/q\mathbb{Z})$ of order $d=2^{\delta}$, and let $b$ be another rational point in $E(\mathbb{Z}/q\mathbb{Z})$ such that $d\,b\neq 0$. Let $\otimes$ denote the multiplication law modulo $b+\langle t\rangle$ defined by Proposition 6. Let $\phi_{\vec{a}}$ (and $^{\mathrm{t}}\phi_{\vec{a}}$ ) denote the linear endomorphism $\vec{x}\mapsto\vec{a}\otimes\vec{x}$ in the coordinate system defined by the basis $(u_{l})_{l\in{\mathbf{Z}}/d{\mathbf{Z}}}$ (and its transposed for the canonical scalar product $\langle\cdot,\,\cdot\rangle$ in this basis).
@@
We note that the elliptic-decisional LWE assumption bears similarities to the decisional Ring-LWE assumption studied by Lyubashevsky, Peikert, and Regev [LPR13], in which matrix products are replaced by polynomial multiplications in the quotient ring $\mathbb{Z}_{q}[x]/\langle\Phi(x)\rangle$ with $\Phi(x)$ a cyclotomic polynomial. When the degree $d$ is a power of two, this construction operates in the cyclotomic field defined by $\Phi_{2d}(x)=x^{d}+1$. Since polynomial multiplication modulo such a polynomial can be implemented using Fast-Fourier-Transform algorithms provided that $2d$ divides q1q-1, this approach achieves quasi-linear time complexity in $d$ for encryption and decryption operations, while also significantly reducing key sizes. This substantial improvement forms the core of cryptosystems currently undergoing standardization [Ava+24, Bai+24].
@@
Nevertheless, it is still possible to make the algorithm in Figure 6 practical using standard techniques from LWE-based cryptography (see, for instance, the Frodo construction [Bos+16]). More precisely, let $\beta$ and $\ell$ be two integer parameters. We can encrypt a $\beta\ell^{2}$-bit message, denoted $(\mu_{i,j})$, by applying the following modifications to the algorithm in Figure 6.
@@
  • Generate secret keys consisting of $\ell$ vectors $\vec{s}_{0}$, $\ldots\vec{s}_{\ell-1}$, associated with public keys $(\vec{a},\vec{w}_{0}$, $\ldots\vec{w}_{\ell-1})$;
@@
Using the same parameters as those employed in standards, for instance $\beta=4$ and $\ell=8$ to encrypt 256-bit data [Bos+16], we observe that this results in timings and data sizes that are not fundamentally larger than in the single-bit version.
More generally, most lattice-based cryptographic schemes roughly follow the following asymptotic guideline, expressed in terms of a security parameter $\lambda$ tending to infinity : $d=O(\lambda\log\lambda)$, $q=O(d)$, and $\chi$ a discrete gaussian distribution with standard deviation $\sigma=O(\sqrt{d})$ [Pei16, Chapter 4]. We can then extract $\beta=O(\log d)$ bits per inner product. Setting furthermore $\ell\simeq O(\sqrt{d}\,/\log d)$ so that we can encrypt on the order of $\lambda$ bits. The total time complexity becomes $O(d^{2}\,/\log d)$ bit-operations, while the data sizes become $O(d^{3/2})$ bits. These complexities are significantly better than the $O(d^{5/2})$ bit-operations of [Bos+16] for same data size.

### v2

@@ 10. Reduction
Under the hypotheses at the beginning of Section 5 and assuming that the base field $\mathbf{K}$} has odd characteristic, we let $b$ be a point on $E$ defined over a separable closure ${\mathbf{K}}_{s}$} of $\mathbf{K}$}. We assume that $db\not=O$. We call $B$ the coset of $b$ under the action of $\langle t\rangle$}. So
@@
We assume that $B$ is left invariant by the absolute Galois group of $\mathbf{K}$}. This is equivalent to the image of $b$ by the isogeny with kernel $\langle t\rangle$} being $\mathbf{K}$}-rational. For every $l$ in ${\mathbf{Z}}/d{\mathbf{Z}}$ we denote by $x_{l}$ the function defined by
@@
There exists a constant $\mathcal{Q}$} such that the following is true. Let $\mathbf{K}$} be a field with odd characteristic. Let $E$ be an elliptic curve over $\mathbf{K}$}. Let $\delta\geqslant 1$ be an integer. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. Let $b$ be a point on $E$ defined over a separable extension of $\mathbf{K}$}. Assume $db\not=O$. Let
@@
This results from the discussion above. To make the procedure clearer, we present the reduction routine as brief pseudo-code in Figure 4. $\square$
@@
Algorithm function Butterfly_Reduce($\vec{F}$)     $\triangleright$ Constants:       _\_ The constants $\vec{{\mathfrak{a}}}\,^{\prime}$ is defined by Equation (8) and $\vec{{\mathfrak{a}}}\,^{1}$ denotes $\vec{{\mathfrak{a}}}\,_{[\,d^{\prime},\ d-1\,]}-\vec{{\mathfrak{a}}}\,_{[\,0,\ d^{\prime}-1\,]}$ .       _\_ The constants $\vec{{\mathfrak{e}}}$, $\vec{{\mathfrak{f}}}$ and $\vec{{\mathfrak{i}}}$ are defined by Equation (33).       _\_ The constant $\vec{{\mathfrak{h}}}$ is defined by Equation (29).       _\_ The constant $\vec{{\mathfrak{l}}}$ is defined by Equation (32), restricted to $l\in[0,\,d^{\prime}-1]$ and
$\mathfrak{l}$$_{\ast}$ denotes the constant ${\mathfrak{l}}_{d^{\prime}}$.
      _\_ The constant $\vec{{\mathfrak{v}}}\,^{\prime}$ denotes $(v^{\prime}_{l}(U^{\prime}))_{l\in{\mathbf{Z}}/d^{\prime}{\mathbf{Z}}}$ .       _\_ The constant $\vec{{\mathfrak{p}}}$ denotes the vector $\vec{{\mathfrak{e}}}\,\diamond\,\vec{{\mathfrak{v}}}\,^{\prime}+\vec{{\mathfrak{f}}}\,\diamond\,(x^{\prime}(U^{\prime}+l\,t^{\prime}))_{l\in{\mathbf{Z}}/d^{\prime}{\mathbf{Z}}}$ , and
$\mathfrak{p}$$_{\ast}$ denotes the constant ${\mathfrak{p}}_{0}-{\mathfrak{f}}_{0}\,x^{\prime}(U^{\prime})-y^{\prime}(U^{\prime})$ .
@@ 11. Multiplication in the residue ring of a fiber
Under the hypotheses and notation of Section 10 we explain how to multiply in the residue ring $\mathbf{L}$} at $B$. We assume that there exists a $\mathbf{K}$}-rational point $R$ on $E$ such that $dR\not=O$. The residue ring $\mathbf{L}$} at $B$ is an algebra of dimension $d$ over $\mathbf{K}$}. We let $(u_{l})_{l\in{\mathbf{Z}}/d{\mathbf{Z}}}$ be the functions on $E$ defined in Section 5. For every $l$ in ${\mathbf{Z}}/d{\mathbf{Z}}$ we let
@@
in $\mathbf{L}$}, given by their coordinates in the basis $\Theta$. We follow [15, Section 4.3.4.]. We first define two functions
@@
This finishes the proof of Proposition 6 below. One can summarize this proof by saying that the equivariant version of Chudnovsky’s multiplication algorithm summarized in [15, Lemma 6] involves convolutions by constant vectors corresponding to evaluation, interpolation and reduction. And the main consequence of the recursive approach presented in Sections 8, 9, and 10 is that these convolution products are achieved in time ${\mathcal{Q}}d\log(d)$ when $d$ is a power of two.
@@
Algorithm function NormalBasis_Multiply($\vec{f},\vec{g}$)     $\triangleright$ The functions $\textsc{Butterfly\_Reduce}_{R}$ and $\textsc{Butterfly\_Reduce}_{B}$ denote the reduction routine defined in Figure 4, where the constants $\mathfrak{h}$} and $\mathfrak{l}$} depend on the points $R\in E({\mathbf{K}})$ and $b\in E({{\mathbf{K}}_{s}})$ as specified in Section 13.         $\vec{\alpha}\leftarrow\textsc{Butterfly\_Evaluate}\,(\vec{f})$   $\vec{\beta}\leftarrow\textsc{Butterfly\_Evaluate}\,(\vec{g})$   $\vec{H}\leftarrow(\vec{f}-\sigma(\vec{f}))\,\diamond\,(\vec{g}-\sigma(\vec{g}))$   $\vec{h}\leftarrow\textsc{Butterfly\_Reduce}_{R}\,(\vec{H})$ $\triangleright$ Reduction modulo $R+{\langle t\rangle}$   $\vec{\gamma}\leftarrow\textsc{Butterfly\_Evaluate}\,(\vec{h})$   $\vec{\delta}\leftarrow\vec{\alpha}\ \diamond\ \vec{\beta}-\vec{\gamma}$   $\vec{k}\leftarrow\textsc{Butterfly\_Interpolate}\,(\vec{\delta})$   $\vec{h}\leftarrow\textsc{Butterfly\_Reduce}_{B}\,(\vec{H})$ $\triangleright$ Reduction modulo $b+{\langle t\rangle}$   return $\vec{h}+\vec{k}$ end function
@@
There exists a constant $\mathcal{Q}$} such that the following is true. Let $\mathbf{K}$} be a field with odd characteristic. Let $E$ be an elliptic curve over $\mathbf{K}$}. Let $\delta\geqslant 1$ be an integer. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. We assume that there exists a point $R$ in $E({\mathbf{K}})$ such that $dR\not=O$. Let $b$ be a point on $E$ defined over a separable extension of $\mathbf{K}$}. Assume $db\not=O$. Let
@@ 12. Implementation and experimental results
A public implementation of the algorithms presented in this work, developed in the computational algebra system Magma [7], is available online [16]. It includes the classical Fast Fourier Transform (FFT) using the Cooley–Tukey algorithm for evaluating or interpolating polynomials in $x$ at $d$-th roots of unity, as well as our routines for elliptic functions at $d$-torsion points, both of which require only $O(d\log d)$ operations in $\mathbb{F}_{q}$. It also covers evaluating or interpolating polynomials in $x$ at $d$-torsion points of an elliptic curve, which requires $O(d\log^{2}d)$ operations using the ECFFT algorithms of [5].
We present in Table 1 the timings for our Magma routines on a standard laptop, working modulo a 64-bit prime. These results are provided for comparison purposes only, given the likely overhead of the Magma interpreter in these highly recursive routines. The relative differences in timings, namely, a constant factor of about $5$ between Cooley-Tukey and the elliptic butterflies, and asymptotically an $O(\log d)$ factor for ECFFT are nonetheless significant in practice. The factor of about 5 reflects, in some sense, the additional technical cost of working with elliptic functions versus polynomials in one variable. Equations (4) are a bit more involved than the classical $x^{a}\cdot x^{b}=x^{a+b}$.
$\log_{2}(d)$ Cooley-Tukey [42] Elliptic Butterflies ECFFT [5]
Evaluate Interpolate Evaluate (Fig. 2) Interpolate (Fig. 3) Enter Exit
8 0.00 s 0.01 s 0.00 s 0.01 s 0.02 s 0.04 s
9 0.01 s 0.01 s 0.02 s 0.01 s 0.05 s 0.11 s
10 0.02 s 0.03 s 0.03 s 0.03 s 0.11 s 0.26 s
11 0.04 s 0.03 s 0.06 s 0.07 s 0.26 s 0.62 s
12 0.03 s 0.03 s 0.13 s 0.15 s 0.60 s 1.42 s
13 0.05 s 0.06 s 0.28 s 0.31 s 1.40 s 3.25 s
14 0.12 s 0.12 s 0.56 s 0.66 s 3.47 s 7.47 s
15 0.25 s 0.25 s 1.17 s 1.38 s 7.65 s 16.9 s
16 0.54 s 0.55 s 2.45 s 2.72 s 17.3 s 38.7 s
@@ 13. Elliptic bases for finite field extensions
We have shown in [15, Theorem 2] how evaluation and interpolation of elliptic functions provide normal bases for finite field extensions having quasi-linear time multiplication. We consider in this section extensions of degree $d=2^{\delta}$, a power of two. We construct normal bases allowing multiplication in time ${\mathcal{Q}}d\log(d)$ for these extensions. We prove the theorem below.
@@
We now consider the case $\nu\leqslant\delta$ and use the constructions in [15]. The integer $d_{q}$ as defined in [15, Definition 1] is either $d^{2}$ or $2d^{2}$. Since $4d^{4}\leqslant q$ we have $d_{q}\leqslant\sqrt{q}$ and we can apply [15, Lemma 9]. There exists an elliptic curve $E$ over $\mathbf{K}$}, a point $t$ of order $d$ in $E({\mathbf{K}})$, a point $R$ in $E({\mathbf{K}})$ such that $dR\not=O$, and a point $b$ in $E({{\mathbf{K}}_{s}})$ such that $db\not=O$ and the conjugate of $b$ by the Frobenius of $E/{\mathbf{K}}$ is $b+t$. We set $B=b+{\langle t\rangle}$. This is a degree $d$ divisor on $E$. It is irreducible over $\mathbf{K}$}. The residue field at $B$ is a degree $d$ extension field of $\mathbf{K}$} which we denote by $\mathbf{L}$}. The $\theta_{l}$ introduced in Equation 34 form a normal basis $\Theta$ of $\mathbf{L}$}. According to Proposition 6 there exists a straight-line program that multiplies to elements in $\mathbf{L}$} given by their coordinates in the basis $\Theta$ at the expense of ${\mathcal{Q}}d\log(d)$ operations in $\mathbf{K}$}. $\square$
@@ 14. Elliptic Goppa codes
The length of the Hasse interval is $4\sqrt{q}$. So there are two consecutive multiples of $d^{2}$ in it. At least one of them is not congruent to $1$ modulo $p$. Call it $m$.
Let c=q+1mc=q+1-m, then $|c|\leq 2\sqrt{q}$ and $m\not\equiv 1\pmod{p}$ is equivalent to $c\not\equiv 0\pmod{p}$. By a theorem of Waterhouse [41], for any integer $c$ with $|c|\leq 2\sqrt{q}$ and $p\nmid c$, there exists an ordinary elliptic curve $E$ over $\mathbb{F}_{q}$ with trace $c$, hence with exactly $m$ rational points (for prime fields this was already proved by Deuring [17]). The group of an elliptic curve over a finite field is either cyclic or a product of two cyclic groups $C_{n_{1}}\times C_{n_{2}}$ with $n_{1}\mid n_{2}$. If $d^{2}$ divides $m$, then $n_{2}$ must be a multiple of $d$. Hence there exists a point of order $d$.
We deduce that there exists an elliptic curve $E$ over $\mathbf{K}$} such that $E({\mathbf{K}})$ contains a point $t$ of order $d$. We set
@@
If the check is failed we deduce that there are errors. We then remind that $C$ is a genus 0 code under its clothes of genus $1$ code. In this context, decoding up to half the minimum distance is achieved thanks to an half-gcd computation as explained in [36]. The time complexity is ${\mathcal{Q}}\log(d)$ times the complexity of multiplying two polynomials of degree $d$ and coefficients in $\mathbf{K}$}. See [40, Chapter 11]. The following theorem summarizes the content of this section
@@ 15. Elliptic LWE cryptography
Our third application concerns the construction of secure cryptographic schemes within quantum computational models, and in particular in the presence of quantum algorithms such as those of Shor [33] and Grover [21]. Among the cryptosystems most extensively studied in this context are those based on computational problems over Euclidean lattices.
@@
Definition 10 (Short-LWE assumptions [29], [30, Lemma 4.4]).
Let $d,q\geq 2$ be integers, let $\chi$ be an error distribution over $\mathbf{Z}$} (typically a discrete Gaussian), and let $\vec{s}\leftarrow\chi^{m}$ be a secret vector. Given pairs $(\mathsf{A},\mathsf{A}\vec{s}+\vec{e})$, where $\mathsf{A}\leftarrow({\mathbf{Z}}/q{\mathbf{Z}})^{m\times d}$ is uniformly random and $\vec{e}\leftarrow\chi^{d}$,
@@
Let $E$ be an elliptic curve defined modulo a prime $q$, with a point $t\in E(\mathbb{Z}/q\mathbb{Z})$ of order $d=2^{\delta}$, and let $b$ be another point in $E(\mathbb{Z}/q\mathbb{Z})$ such that $db\neq O$. Let $\otimes$ denote the multiplication law modulo $b+\langle t\rangle$ defined in Proposition 6. Finally, let $\chi$ be an error distribution over $\mathbb{Z}$ (typically a discrete Gaussian), and let $\vec{s}\leftarrow\chi^{d}$ be a secret vector.
@@
Transposing the Regev construction in this setting yields a secure encryption scheme in the chosen-plaintext attack (CPA) model.
@@
The correctness follows from the equation $p=\langle\vec{r}\,,\vec{e}\rangle-\langle\vec{e}_{1}\,,\vec{s}\rangle+e_{2}+\mu\,\lfloor q/2\rceil\,.$ Since the error term remains small for appropriate parameter choices, decryption succeeds with high probability. CPA-security follows directly from the decisional LWE assumption. The vector $\vec{w}$ is indistinguishable from a random vector, which makes $\vec{c}_{1}$ and $\langle\vec{r},\,\vec{w}\rangle+e_{2}$ two LWE instances, each indistinguishable from uniform, thereby perfectly hiding the message.
In terms of complexity, applying $\phi_{\vec{a}}$ requires $\mathcal{Q}\,d\log(d)$ operations in $\mathbb{Z}/q\mathbb{Z}$, using the algorithms described in Section 10. By Tellegen’s principle, applying the transpose map $^{\mathrm{t}}\phi_{\vec{a}}$ has the same asymptotic complexity [25, 8]. More generally, all steps can be carried out within this time bound. The sizes of the keys and ciphertexts are $O(d)$ bits.
Parameters Let $E$ be an elliptic curve defined modulo a prime $q$ with a point $t\in E(\mathbb{Z}/q\mathbb{Z})$ of order $d=2^{\delta}$, and let $b$ be another rational point in $E(\mathbb{Z}/q\mathbb{Z})$ such that $d\,b\neq O$. Let $\otimes$ denote the multiplication law modulo $b+\langle t\rangle$ defined by Proposition 6. Let $\phi_{\vec{a}}$ (and $^{\mathrm{t}}\phi_{\vec{a}}$ ) denote the linear endomorphism $\vec{x}\mapsto\vec{a}\otimes\vec{x}$ in the coordinate system defined by the basis $(u_{l})_{l\in{\mathbf{Z}}/d{\mathbf{Z}}}$ (and its transposed for the canonical scalar product $\langle\cdot,\,\cdot\rangle$ in this basis).
@@
We note that the elliptic-decisional LWE assumption bears similarities to the decisional Ring-LWE assumption studied by Lyubashevsky, Peikert, and Regev [26], in which matrix products are replaced by polynomial multiplications in the quotient ring $\mathbb{Z}_{q}[x]/\langle\Phi(x)\rangle$ with $\Phi(x)$ a cyclotomic polynomial. When the degree $d$ is a power of two, this construction operates in the cyclotomic field defined by $\Phi_{2d}(x)=x^{d}+1$. Since polynomial multiplication modulo such a polynomial can be implemented using Fast-Fourier-Transform algorithms provided that $2d$ divides q1q-1, this approach achieves quasi-linear time complexity in $d$ for encryption and decryption operations, while also significantly reducing key sizes. This substantial improvement forms the core of cryptosystems currently undergoing standardization [1, 2].
@@
Nevertheless, it is still possible to make the algorithm in Figure 6 practical using standard techniques from LWE-based cryptography (see, for instance, the Frodo construction [6]). More precisely, let $\beta$ and $\ell$ be two integer parameters. We can encrypt a $\beta\ell^{2}$-bit message, denoted $(\mu_{i,j})$, by applying the following modifications to the algorithm in Figure 6.
@@
  • Generate secret keys consisting of $\ell$ vectors $\vec{s}_{0}$, $\ldots,\,\vec{s}_{\ell-1}$, associated with public keys $(\vec{a},\vec{w}_{0}$, $\ldots,\,\vec{w}_{\ell-1})$;
@@
Using the same parameters as those employed in standards, for instance $\beta=4$ and $\ell=8$ to encrypt 256-bit data [6], we observe that this results in timings and data sizes that are not fundamentally larger than in the single-bit version.
More generally, most lattice-based cryptographic schemes roughly follow the following asymptotic guideline, expressed in terms of a security parameter $\lambda$ tending to infinity : $d=O(\lambda\log\lambda)$, $q=O(d)$, and $\chi$ a discrete gaussian distribution with standard deviation $\sigma=O(\sqrt{d})$ [27, Chapter 4]. We can then extract $\beta=O(\log d)$ bits per inner product. Setting furthermore $\ell\simeq O(\sqrt{d}\,/\log d)$ so that we can encrypt on the order of $\lambda$ bits. The total time complexity becomes $O(d^{2}\,/\log d)$ bit-operations, while the data sizes become $O(d^{3/2})$ bits. These complexities are significantly better than the $O(d^{5/2})$ bit-operations of [6] for same data size.